# Incorporating Task Progress Knowledge for Subgoal Generation in Robotic Manipulation through Image Edits

### v1

@@ Abstract
Understanding the progress of a task allows humans to not only track what has been done but also to better plan for future goals. We demonstrate TaKSIE, a novel framework that incorporates task progress knowledge into visual subgoal generation for robotic manipulation tasks. We jointly train a recurrent network with a latent diffusion model to generate the next visual subgoal based on the robot’s current observation and the input language command. At execution time, the robot leverages a visual progress representation to monitor the task progress and adaptively samples the next visual subgoal from the model to guide the manipulation policy. We train and validate our model in simulated and real-world robotic tasks, achieving state-of-the-art performance on the CALVIN manipulation benchmark. We find that the inclusion of task progress knowledge can improve the robustness of trained policy for different initial robot poses or various movement speeds during demonstrations. The project website can be found at https://live-robotics-uva.github.io/TaKSIE/.
@@ 1 Introduction
Teaching robots to perform complex tasks usually involves turning a task into a sequence of subgoals, several steps towards the overall task goal. Several hierarchical methods have been proposed to generate subgoals and have low-level policy to learn to reach the generated subgoals, e.g., task and motion planning [8], options in reinforcement learning [35], and learning hierarchical policies [42, 30].
Effectively generating subgoals requires the robot to understand how a task may progress and the key steps for completing a task. Without this knowledge, the robot may over- or under-generate subgoals. Over-generating subgoals may lead to inefficiency in plan execution because this introduces unnecessary intermediate steps for the robot to achieve. Under-generating subgoals may hurt task performance as each subgoal is still hard for the robot to achieve. Furthermore, if the robot does not understand the task progress, it may synthesize inconsistent subgoals, especially when the solution distribution is multimodal. For example, for task “pick up a cup”, if there are multiple cups, the robot may generate the first subgoal to move towards one cup, but have the second subgoal that makes the robot move to another cup at midway. Only when having the task progress knowledge, the robot can generate the subgoals that represent the key steps of a task as shown in Fig. 1.
@@ 2.1 Learning Goal-Conditioned Policies
Goal-conditioned policies enable robots to generate actions for a specified goal, either represented in a natural language sentence [21, 22] or an image depicting the desired state [28]. Most approaches use a pretrain text or image encoder to turn the language or visual goals into embeddings to condition the policy network to generate actions. While language goals are more expressive, it is harder to acquire annotated demonstration data. On the other hand, the visual goals can only represent one specific goal configuration, we can leverage large amount of unannotated trajectories to train policy. In TaKSIE, our subgoal generation process effectively turns a language-goal-conditioned problem into a visual-goal-conditioned problem so it is possible to leverage unannotated trajectories to solve the manipulation tasks specified in natural language.
@@ 2.2 Planning with Diffusion Models
Diffusion models have demonstrated the ability to effectively approximate and re-generate data distributions such as in image generation [12, 41] and robotic action generation [14, 1, 40, 4]. Several diffusion based policy models have been proposed, for example, diffuser [14] refine policies from noise, decision diffuser [1] integrates constraints and skills, and LCD [39] conditions latent plans on language for task execution. Different from these policy models that generate actions, our approach generates visual subgoals with diffusion models to enhance policy guidance. The closest model to us is SuSIE [3] which also leverages the diffusion model for visual subgoal generation. Unlike SuSIE which uses fixed intervals to generate subgoals, we focus on incorporating and tracking task progress in the subgoal generation process. Other than image-based subgoals, other works explore generating short videos to guide the policy [2, 7, 17]. However, generating full video sequences usually requires much more computing resources.
@@ 2.3 Task Understanding for Robotic Tasks
To understand what the next subgoals can be, the robot needs to learn the relevant key steps for a task. Such knowledge has been encoded as task plans in predefined domains [9, 15]. Another way to represent task progress is using visual keyframes [25, 16]. Recently, large language models (LLMs) have demonstrated the ability to generate task plans or control programs for a variety of robotic tasks [31, 18]. However, LLM-based approaches require converting a state into textual descriptions to synthesize the prompts and assume a set of API calls exists. To generate visual subgoals that are useful for a visual-goal-conditioned policy, we need to consider task knowledge in the subgoal generation process. Several prior works have explored using human videos to help robot learning[32, 37, 29] as they implicitly encode task knowledge. To scale this up, recent works like R3M [24], LIV [19], and VC-1 [20] train on diverse and large-scale human video data to learn useful visual task representations for robotic tasks. Our approach is a keyframe-based approach but leverages these visual representations to measure task progress and enhance the consistency in generation using a progress encoder.
@@ 3.3 Incorporating Task Progress
To consider task progress in subgoal generation, we need to first have a measure of whether a state is making progress toward completing a task. Then we can use this measure to 1) identify the key states for a task and train the subgoal generation using the identified key states and 2) evaluate whether a subgoal is reached. Recent visual representations pretrained using temporal contrastive learning on ego-centric videos, such as R3M [24] and LIV [19], implicitly encode the task progress knowledge. They are usually trained by minimizing the distance between temporally close frames while maximizing distant or unrelated ones, which can capture the essential features for sequential task progression. In TaKSIE, we utilize this pretrained visual representation $\mathcal{M}$ to measure the task progress.
@@ 3.3.1 Ground-truth Subgoal Selection
This definition means that these are the points we need to reach first, then to make the obvious progress. Fig. 3 shows an example of measured task progress using this distance. The two selected subgoals for the task “open the drawer” are shown in the right of Fig. 3 indicating that the robot needs to first move above the drawer handle and then insert its gripper into the handle.
@@ 3.3.2 Progress Encoder
Selecting the key steps as the ground-truth subgoals for training the subgoal generator only help instruct the robot to generate subgoals that are close to key steps. However, many tasks can be solved in different ways, i.e., multimodal solutions. For example, to slide the door to the right, the robot can choose to grab the handle to move the door to the right or put its gripper to the left of the handle to push it to the right. The generation of two consecutive subgoals may still be inconsistent so the robot will switch between modes and fail to finish the task. To enforce consistency in subgoal generation, we create an external memory, a progress encoder, to keep track of the subgoals achieved so far and guide the next subgoal. The progress encoder is a recurrent network that takes the image of the current state $s_{t}$ and the last hidden state $h_{t-1}$ of the progress encoder to update its hidden state. The output hidden state $h_{t}$ is the progress embedding that is used in the subgoal generator.
@@ 3.3.3 Progress Evaluator
Again, we use the pretrained time-contrastive visual representations to measure if the current state and the generated subgoal are in similar progress. We use cosine similarity as the similarity measurement.
If the similarity reaches $\delta$ or the step for the current subgoal is larger than $\lambda$, it will move forward to the next subgoal. This process ensures continuous progress toward the achievement of the overall goal.
@@ 3.4.1 Subgoal Generator
The subgoal generator is a conditional generative model. We implement this generator by combining the 2D latent diffusion model [27] and ControlNet as in  [41] to ensure the generation follows the conditions:
@@
The image of the current state $s_{t}$ is input to the ControlNet branch so the generated image includes objects in the current state. The language goal is encoded by a pretrained text encoder, e.g., the CLIP text encoder. This embedding of the goal is fused with the progress embedding $h_{t}$ and the diffusion time step to condition on each level of the U-Net. Conditioning on goal and progress on each level allows us to generate images that match the next subgoal.
@@ 3.5 Training Pipeline
We first fine-tune the diffusion model with the selected ground-truth subgoal images. This fine-tuning process allows us to generate images that match the style of the environment. We then initialize ControlNet with the weights obtained from the fine-tuned diffusion model and then co-train it with the progress encoder. This co-training process enables us to jointly learn to track the task progress and accurately predict the next visual subgoal in the sequence. To ensure the successful prediction of the next subgoal, we adopt the MSE loss, $\mathcal{L}_{gen}$, to compare the generated subgoal, $f_{gen}$ and the ground-truth subgoal, $f_{gt}$:
@@
We train the policy network with goal-conditioned behavior cloning (GCBC). To increase the variety of the initial state and the goal state observed by the goal-conditioned policy, we randomly sample subtrajectory from the human demonstration dataset $\mathcal{D}$ using window sizes ranging from $k_{\min}$ to $k_{\max}$. To better reach the final goal, we repeat the final state $k_{\delta}$ times. In our implementation, we use a diffusion policy based GCBC (D-GCBC) and thus use a similar loss function as Eq. 2 but we replace image $f$ with action $\mathbf{a}$.
@@ 4.1 Dataset and Evaluation Environment
To show our proposed framework can be integrated with the real robot. We designed 28 tasks that resemble the tasks in CALVIN. For example, opening the drawer, turning on the red LED, placing the red block in the box, etc. For each task, we used an Oculus VR controller to teleoperate the robot to collect 50 trajectories to use as our training set. We record the trajectories at 30Hz but downsample to 15Hz for training. Each trajectory includes a language command for the task, a sequence of still images from an over-the-shoulder camera, and a sequence of end-effector pose and gripper state corresponding to the image observation. Unlike the trajectories in CAVLIN which are of similar length, the real-world trajectories have various trajectory lengths ranging from 10 to 104. At the evaluation phase, we randomized our environment configuration and tested each task 5 times.
@@ 4.2.1 Baselines
  • SuSIE [3]: Similar to our approach, SuSIE employs diffusion models to generate subgoals that guide the execution of low-level policies, but it was trained and evaluated at fixed intervals. We retrain SuSIE using the implementation and configuration provided in the original repository.
@@ 4.2.2 Implementation Details
We use the selected visual representation to calculate the distance of each frame to the final frame. Then we smooth the distance curve using a smooth factor 0.167, and set $\delta_{1}=0.001$, $\delta_{2}=-0.001$, and minimum subgoal interval $d=7$ to avoid the impact of outliers. On average, in CALVIN, we select 2.8 ground-truth subgoals, including the final goal, with an interval of 20; in the real robot experiment, we select 1.6 ground-truth subgoals with an interval of 23.
@@
In all experiments, we use the pretrained weights from [26] to initialize our diffusion model. However, training a high-quality generative model requires a fair amount of images. Instead of only training with subgoal images, we first fine-tune the unconditional diffusion model using all frames in the recorded trajectories. On average, there are 64 images per trajectory in CALVIN and 32 images per trajectory for the real robot dataset. The unconditioned diffusion model is trained for 100k steps with batch size 256. This is a key step to generate realistic images. After this, we can effectively use the subgoal images to train the conditional model to follow progress. We found that training the diffusion model together with the ControlNet branch yielded a better image quality than keeping the diffusion model part frozen. We use the weights of the trained unconditional diffusion model to initialize the ControlNet part. Finally, we trained the diffusion model and ControlNet together with batch size 256 for 300k training steps. This training procedure allows us to effectively use the limited amount of demonstration data to generate high-quality subgoal images.
@@
We use a diffusion model based goal-conditioned policy (D-GCBC) [36] as our policy network. We stack the current observation and the goal image, encode them with ResNet-50, and then use this embedding to condition a 3 256-unit layers MLP model. Following [4, 3], we predict four action sequences for temporal consistency and do a dimension-wise mean of them. In CALVIN, we set $k_{min}=8$ and $k_{max}=20$ as our minimum and maximum ground-truth subgoal interval is 8 and 32. In real-world experiments, we set $k_{min}=4$ and $k_{max}=20$. Both models are trained in 400k steps with batch size 256.
@@ 4.2.3 Policy Rollout
At inference, our subgoal generator uses DDIM [33] with 50 time steps and with text and image classifier-free guidance scales set to 2.5. Our progress evaluator sets the cosine similarity threshold $\delta$ to be 0.96. The maximum number of rollout steps $\lambda$ for a subgoal image is 20.
@@ 4.3.1 Quality of Generated Subgoals
To assess the quality of the generated subgoals, we compare our task success rate against the policy conditioned on the ground-truth (GT) goal images. The condition that only uses the GT final goal shows the performance of good image quality of the goal but does not consider the task progress so only one goal is used. The condition that uses the GT subgoal images uses our progress evaluation to determine whether to move to the next subgoal. This represents the performance of subgoals when the subgoals have good image quality and good understanding of task progress. During rollouts, we use the same policy network and parameters for these conditions. We run this evaluation on the CALVIN validation set which has 34 tasks, and around 30 rollouts per task. Tab. 1 showed that the subgoals generated by TaKSIE can lead to a similar success rate as using GT subgoals, indicating that our generated subgoals are as good as the ones selected from the human demonstrations.
@@ 4.3.5 Ablations: Visual Representations
The bottom two rows in Tab. 2 showed the results of different visual representations. R3M performs best. Compared to CLIP, R3M encodes progress better as it is trained with time-contrastive learning which distinguishes frames more effectively. Compared with LIV which is pretrained on the EPIC-KITCHENS [6], R3M is trained on Ego4D [10] which is a much larger ego-centric video dataset so it shows better generalization to out of domain videos.

### v2

@@ Abstract
Understanding the progress of a task allows humans to not only track what has been done but also to better plan for future goals. We demonstrate TaKSIE, a novel framework that incorporates task progress knowledge into visual subgoal generation for robotic manipulation tasks. We jointly train a recurrent network with a latent diffusion model to generate the next visual subgoal based on the robot’s current observation and the input language command. At execution time, the robot leverages a visual progress representation to monitor the task progress and adaptively samples the next visual subgoal from the model to guide the manipulation policy. We train and validate our model in simulated and real-world robotic tasks, achieving state-of-the-art performance on the CALVIN manipulation benchmark. We find that the inclusion of task progress knowledge can improve the robustness of trained policy for different initial robot poses or various movement speeds during demonstrations. The project page is available at https://live-robotics-uva.github.io/TaKSIE/.
@@ 1 Introduction
Teaching robots to perform complex tasks usually involves turning a task into a sequence of subgoals, several steps towards the overall task goal. Several hierarchical methods have been proposed to generate subgoals and have low-level policies to learn to reach the generated subgoals, e.g., task and motion planning [9], options in reinforcement learning [35], and learning hierarchical policies [42, 30].
Effectively generating subgoals requires the robot to understand how a task may progress and the key steps for completing a task. Without this knowledge, the robot may over- or under-generate subgoals. Over-generating subgoals may lead to inefficiency in plan execution because this introduces unnecessary intermediate steps for the robot to achieve. Under-generating subgoals may hurt task performance as each subgoal is still hard for the robot to achieve. Furthermore, if the robot does not understand the task progress, it may synthesize inconsistent subgoals, especially when the solution distribution is multimodal. For example, for the task “pick up a cup”, if there are multiple cups, the robot may generate the first subgoal to move towards one cup, but have the second subgoal that makes the robot move to another cup at midway. Only when having the task progress knowledge, the robot can generate the subgoals that represent the key steps of a task as in Fig. 1.
@@ 2.1 Learning Goal-Conditioned Policies
Goal-conditioned policies enable robots to generate actions for a specified goal, either represented in a natural language sentence [22, 23] or an image depicting the desired state [28]. Most approaches use a pretrain text or image encoder to turn language or visual goals into embeddings to condition the policy network to generate actions. While language goals are more expressive, it is harder to acquire annotated demonstration data. On the other hand, visual goals can only represent one specific goal configuration, and we can leverage a large amount of unannotated trajectories to train policy. In TaKSIE, our subgoal generation process effectively turns a language-goal-conditioned problem into a visual-goal-conditioned problem, so it is possible to leverage unannotated trajectories to solve the manipulation tasks specified in natural language.
@@ 2.2 Planning with Diffusion Models
Diffusion models have demonstrated the ability to effectively approximate and re-generate data distributions such as in image generation [13, 41] and robotic action generation [15, 1, 40, 5]. Several diffusion-based policy models have been proposed, for example, diffuser [15] refine policies from noise, decision diffuser [1] integrates constraints and skills, and LCD [39] conditions latent plans on language for task execution. Different from these policy models that generate actions, our approach generates visual subgoals with diffusion models to enhance policy guidance. The model closest to us is SuSIE [3] which also leverages the diffusion model for the generation of visual subgoals. Unlike SuSIE which uses fixed intervals to generate subgoals, we focus on incorporating and tracking task progress in the subgoal generation process. Other than image-based subgoals, other works explore generating short videos to guide the policy [2, 8, 18]. However, generating full video sequences usually requires much more computing resources.
@@ 2.3 Task Understanding for Robotic Tasks
To understand what the next subgoals can be, the robot needs to learn the relevant key steps for a task. This knowledge has been encoded as task plans in predefined domains [10, 16]. Another way to represent task progress is using visual keyframes [26, 17]. Recently, large language models (LLMs) have demonstrated the ability to generate task plans or control programs for a variety of robotic tasks [31, 19]. However, LLM-based approaches require converting a state into textual descriptions to synthesize the prompts and assume that a set of API calls exists. To generate visual subgoals that are useful for a visual-goal-conditioned policy, we need to consider task knowledge in the subgoal generation process. Several previous works have explored the use of human videos to help robot learning[32, 37, 29] as they implicitly encode task knowledge. To scale this up, recent works like R3M [25], LIV [20], and VC-1 [21] train on diverse and large-scale human video data to learn useful visual task representations for robotic tasks. Our approach is keyframe-based but leverages these visual representations to measure task progress and enhance the consistency in generation using a progress encoder.
@@ 3.3 Incorporating Task Progress
To consider task progress in subgoal generation, we need to first have a measure of whether a state is making progress toward completing a task. Then we can use this measure to 1) identify the key states for a task and train the subgoal generation using the identified key states and 2) evaluate whether a subgoal is reached. Recent visual representations pretrained using time contrastive learning on ego-centric videos, such as R3M [25] and LIV [20], implicitly encode the task progress knowledge. They are usually trained by minimizing the distance between temporally close frames while maximizing distant or unrelated ones, which can capture the essential features for sequential task progression. In TaKSIE, we utilize this pretrained visual representation $\mathcal{M}$ to measure the task progress.
@@ 3.3.1 Ground-truth Subgoal Selection
This definition means that these are the points we need to reach first, then to make obvious progress. Fig. 3 shows an example of measured task progress using this distance. The two selected subgoals for the task “open the drawer” are shown in the right of Fig. 3 indicating that the robot needs to first move above the drawer handle and then insert its gripper into the handle.
@@ 3.3.2 Progress Encoder
Selecting the key steps as the ground-truth subgoals for training the subgoal generator only helps instruct the robot to generate subgoals that are close to key steps. However, many tasks can be solved in different ways, i.e., multimodal solutions. For example, to slide the door to the right, the robot can choose to grab the handle to move the door to the right or put its gripper to the left of the handle to push it to the right. The generation of two consecutive subgoals may still be inconsistent so the robot will switch between modes and fail to finish the task. To enforce consistency in subgoal generation, we create an external memory, a progress encoder, to keep track of the subgoals achieved so far and guide the next subgoal. The progress encoder is a recurrent network that takes the image of the current state $s_{t}$ and the last hidden state $h_{t-1}$ of the progress encoder to update its hidden state. The output hidden state $h_{t}$ is the progress embedding that is used in the subgoal generator.
@@ 3.3.3 Progress Evaluator
Again, we use the pretrained time-contrastive visual representations to measure if the current state and the generated subgoal are in similar progress. We use cosine similarity as the similarity measurement. If the similarity reaches $\delta$ or the step for the current subgoal is larger than $\lambda$, it will move forward to the next subgoal. This process ensures continuous progress toward the achievement of the overall goal.
@@ 3.4.1 Subgoal Generator
The subgoal generator is a conditional generative model. We implement this generator using a 2D image-conditioned diffusion model, allowing flexibility in the choice of the model, e.g., ControlNet [41] or InstructPix2Pix [4]. A detailed analysis of these choices can be found in Appx. D. The diffusion model conditions on image observations to ensure the generation follows the conditions:
@@
The image of the current state $s_{t}$ is used as input to the conditioned generative model to ensure the generated subgoal image includes objects in the current state.
The language goal is encoded by a pretrained text encoder, e.g., the CLIP text encoder. This embedding of the goal is fused with the progress embedding $h_{t}$ and the diffusion time step to condition on each level of the U-Net. Conditioning on goal and progress on each level allows us to generate images that match the next subgoal.
@@ 3.5 Training Pipeline
We first fine-tune the diffusion model with the selected ground-truth subgoal images. This fine-tuning process allows us to generate images that match the style of the environment. We then initialize the conditioned generative model with weights from the fine-tuned diffusion model. For ControlNet, we copy the weights to the branch and co-train with the U-Net branch. For InstructPix2Pix, we apply the weights to the full model and train. This co-training process enables us to jointly learn to track the task progress and accurately predict the next visual subgoal in the sequence. To ensure the successful prediction of the next subgoal, we adopt the MSE loss, $\mathcal{L}_{gen}$, to compare the generated subgoal, $f_{gen}$ and the ground-truth subgoal, $f_{gt}$:
@@
We train the policy network with goal-conditioned behavior cloning (GCBC). To increase the variety of the initial state and the goal state observed by the goal-conditioned policy, we randomly sample subtrajectories from the human demonstration dataset $\mathcal{D}$ using window sizes ranging from $k_{\min}$ to $k_{\max}$. To better reach the final goal, we repeat the final state $k_{\delta}$ times. In our implementation, we use a diffusion policy based GCBC (D-GCBC) and thus use a similar loss function as Eq. 2 but we replace image $f$ with action $\mathbf{a}$.
@@ 4.1 Dataset and Evaluation Environment
To show our proposed framework can be integrated with the real robot, Franka Research 3. We designed 28 tasks that resemble the tasks in CALVIN. For example, opening the drawer, turning on the red LED, placing the red block in the box, etc. For each task, we used an Oculus VR controller to teleoperate the robot to collect 50 trajectories to use as our training set. We record the trajectories at 30Hz but downsample to 15Hz for training. Each trajectory includes a language command for the task, a sequence of still images from an over-the-shoulder camera, and a sequence of end-effector pose and gripper state corresponding to the image observation. Unlike the trajectories in CAVLIN which are of similar length, the real-world trajectories have various trajectory lengths ranging from 10 to 104. At the evaluation phase, we randomized our environment configuration and tested each task 5 times.
@@ 4.2.1 Baselines
  • SuSIE [3]: Similar to TaKSIE, SuSIE employs diffusion models to generate subgoals that guide the low-level policy, but it was trained and evaluated at fixed intervals. We retrain SuSIE using the implementation and configuration provided in the original repository.
@@ 4.2.2 Implementation Details
We use the selected visual representation to calculate the distance of each frame to the final frame. Then we smooth the distance curve using a smooth factor 0.167, and set $\delta_{1}=0$, $\delta_{2}=0$, and minimum subgoal interval $d=7$ to avoid the impact of outliers. On average, in CALVIN, we select 2.74 ground-truth subgoals, including the final goal, with an interval of 16.4; in the real robot experiment, we select 1.6 ground-truth subgoals with an interval of 23.
@@
In all experiments, we use the pretrained weights from [27] to initialize our diffusion model. However, training a high-quality generative model requires a fair amount of images. Instead of only training with subgoal images, we first fine-tune the unconditional diffusion model using all frames in the recorded trajectories. On average, there are 64 images per trajectory in CALVIN and 32 images per trajectory for the real robot dataset. The unconditioned diffusion model is trained for 100k steps with batch size 256. This is a key step to generate realistic images. After this, we can effectively use the subgoal images to train the conditional model to follow progress. The original ControlNet freezes the U-Net branch and trains only the ControlNet branch. We found that training the diffusion model together with the ControlNet branch yielded a better image quality than keeping the diffusion model part frozen. We use the weights of the trained unconditional diffusion model to initialize the ControlNet part. Finally, we trained InstructPix2Pix or ControlNet with batch size 256 for 300k training steps. This training procedure allows us to effectively use the limited amount of demonstration data to generate high-quality subgoal images.
@@
We use a diffusion model based goal-conditioned policy (D-GCBC) [36] as our policy network. We stack the current observation and the goal image, encode them with ResNet-50, and then use this embedding to condition a 3 256-unit layers MLP model. Following [5, 3], we predict four action sequences for temporal consistency and do a dimension-wise mean of them. In CALVIN, we set $k_{\textit{min}}=8$ and $k_{\textit{max}}=20$ as our minimum and maximum ground-truth subgoal interval is 8 and 32. In real-world experiments, we set $k_{\textit{min}}=4$ and $k_{\textit{max}}=20$. Both models are trained in 400k steps with batch size 256.
@@ 4.2.3 Policy Rollout
At inference, our subgoal generator uses DDIM [33] with 50 time steps and with text and image classifier-free guidance scales set to 2.5. Our progress evaluator sets the cosine similarity threshold $\delta$ to 0.96. The maximum number of rollout steps $\lambda$ for a subgoal image is 20.
@@ 4.3.1 Quality of Generated Subgoals
To assess the quality of the generated subgoals, we compare our task success rate against the policy conditioned on the ground-truth (GT) goal images. The condition that only uses the GT final goal shows the performance of good image quality of the goal but does not consider the task progress so only one goal is used. The condition that uses the GT subgoal images uses our progress evaluation to determine whether to move to the next subgoal. This represents the performance of subgoals when the subgoals have good image quality and a good understanding of task progress. During rollouts, we use the same policy network and parameters for these conditions. We run this evaluation on the CALVIN validation set which has 34 tasks, and around 30 rollouts per task. Tab. 1 showed that the subgoals generated by TaKSIE can lead to a similar success rate as using GT subgoals, indicating that our generated subgoals are as good as the ones selected from the human demonstrations.
@@ 4.3.5 Ablations: Visual Representations
The bottom two rows in Tab. 2 show the results of different visual representations. R3M performs best. Compared to CLIP, R3M encodes progress better as it is trained with time-contrastive learning which distinguishes frames more effectively. Compared with LIV which is pretrained on the EPIC-KITCHENS [7], R3M is trained on Ego4D [11] which is a much larger ego-centric video dataset so it shows better generalization to out of domain videos.